Using a Truth table, determine the value of the compound proposition ((𝑝 ∨ 𝑞) ∧ (¬𝑝 ∨ 𝑟)) → (𝑞 ∨ 𝑟).


1
Expert's answer
2021-05-02T07:53:51-0400

First, remember the truth tables for basic logical operations :




Before constructing a truth table for the specified expression, we will define a sequence of logical operations. We will denote by numbers from above operations.



"\left(\left(p\lor^{(1)} q\right)\land^{(4)}\left(\neg p\lor^{(2)} r\right)\right)\longrightarrow^{(5)}\left(q\lor^{(3)} r\right)"

Now we can construct a truth table for the specified expression.





Q.E.D.